Given a standard normal distribution with mean =101 and standard deviation = 43, find P(X>145)


1
Expert's answer
2022-02-03T17:50:41-0500

"X\sim N(101, 43^2)"



"P(X>145)=1-P(X\leq 145)""=1-P(Z\leq \dfrac{145-101}{43})=1-P(Z\leq1.0232)=1-0.84689=0.1531"